In the number 7^1000, what is the digit in the units place?
Answer by oscargut(891)

7^1=7
7^2=49
7^3=343
7^4 ends in 1
7^5 ends in 7
7^6 ends in 9
7^7 ends in 3
7^8 ends in 1
7^9 ends in 7
then 7^n ends in 1 if only if 4 is a divisor of n
for example 7^8, 7^44, 7^100 and 7^1000 ends in 1
Answer:7^1000 ends in 1